### v1

# RealEngine: Simulating Autonomous Driving in Realistic Context

Abstract
Driving simulation plays a crucial role in developing reliable driving agents by providing controlled, evaluative environments. To enable meaningful assessments, a high-quality driving simulator must satisfy several key requirements: multi-modal sensing capabilities (e.g., camera and LiDAR) with realistic scene rendering to minimize observational discrepancies; closed-loop evaluation to support free-form trajectory behaviors; highly diverse traffic scenarios for thorough evaluation; multi-agent cooperation to capture interaction dynamics; and high computational efficiency to ensure affordability and scalability. However, existing simulators and benchmarks fail to comprehensively meet these fundamental criteria. To bridge this gap, this paper introduces RealEngine, a novel driving simulation framework that holistically integrates 3D scene reconstruction and novel view synthesis techniques to achieve realistic and flexible closed-loop simulation in the driving context. By leveraging real-world multi-modal sensor data, RealEngine reconstructs background scenes and foreground traffic participants separately, allowing for highly diverse and realistic traffic scenarios through flexible scene composition. This synergistic fusion of scene reconstruction and view synthesis enables photorealistic rendering across multiple sensor modalities, ensuring both perceptual fidelity and geometric accuracy. Building upon this environment, RealEngine supports three essential driving simulation categories: non-reactive simulation, safety testing, and multi-agent interaction, collectively forming a reliable and comprehensive benchmark for evaluating the real-world performance of driving agents.
1 Introduction
Autonomous driving (AD) methods yin2021center ; li2022homogeneous ; ye2023fusionad ; hu2023planning ; wen2023dilu ; fu2024drive have advanced rapidly, largely due to the introduction of diverse driving datasets nuplan ; caesar2020nuscenes ; montali2024waymo and simulators dosovitskiy2017carla . These resources facilitate research by supporting model training, testing, and evaluation across a variety of virtual and real driving environments, utilizing a wealth of multimodal sensor, including cameras and LiDAR.
However, existing AD datasets and simulators have fundamental limitations that prevent them from fully capturing real-world driving scenarios and challenges, diminishing their credibility and usefulness in practical applications. For instance, real driving data typically provides only pre-existing driving trajectories nuplan ; caesar2020nuscenes , allowing for open-loop evaluation without immediate feedback or interaction with the trajectories planned by driving agents. This results in a significant discrepancy between simulated and actual driving behavior. The static nature of recorded datasets, where other vehicles do not react to the actions of the ego vehicle, creates a scenario fundamentally different from real-world situations. Although the closed-loop simulator CARLA dosovitskiy2017carla addresses this issue by providing real-time feedback to driving agents, it relies on manual 3D modeling and graphics engines, which lacks realism and exhibits a substantial appearance gap from actual driving scenarios. Consequently, models trained in this environment may struggle to handle real-world driving conditions effectively. Further, previous benchmarks primarily focus on non-collision and normal driving scenarios, which limits the models’ ability to address unseen risks encountered in the real world.
To advance the field, we introduce RealEngine, a pioneering autonomous driving simulation platform capable of rendering realistic multimodal sensor data efficiently and supporting closed-loop simulation. It is distinguished by the following features: (i) Realistic scene rendering, closely resembling the real world to minimize domain discrepancies between simulated and actual driving environments, allowing both camera images and LiDAR point clouds; (ii) Closed-loop simulation, enabling driving along free-form trajectories planned by agents while providing corresponding feedback; (iii) Supporting diverse scenarios, including a wide range of hazardous driving situations to facilitate comprehensive safety-critical evaluations of driving agents; (iv) Multi-agent co-operation and interaction, closely approximating various real-world conditions in terms of driving dynamics and scene complexity. As summarized in Table 1, no existing benchmarks or simulators meet these fundamental requirements simultaneously, as well as a spectrum of driving simulation focused features.
Our RealEngine is founded on the innovative concept of seamlessly integrating scene reconstruction with the composition of traffic participants. We reconstruct the background scene and foreground traffic participants separately with the corresponding real-world data. This approach simultaneously addresses key limitations of existing scene reconstruction methods chen2023periodic ; wu2023mars ; yang2023emernerf ; yan2024street ; chen2024omnire ; turki2023suds : (i) Unable to acquire occluded regions within a scene; (ii) Inferior performance in synthesizing novel viewpoints of traffic participants; (iii) Limited choice of possible traffic participants. Our decomposed design naturally supports flexible scene editing while enabling the concurrent operation of multiple driving agents, as required for realistic driving simulations. By merging the background scene with diverse traffic participants, we can efficiently simulate a wide range of high-quality, unique driving scenarios tailored to any specific requirements. We generate a diverse range of driving scenarios and simulate three driving categories: non-reactive, safety test, and multi-agent interaction. This offers a high quality, reliable, flexible, and comprehensive benchmark for assessing the real-world performance of driving agents.
Table 1: Comparison of various datasets, generative models, world models, and simulators in terms of interactivity, fidelity, diversity, and efficiency. DATA. represents dataset, GEN. represents generative model, W.M. represents world model, SIM. represents simulator.
Type Name Interactivity Fidelity Diversity Efficiency
Uncontrollable
closed-loop
Controllable
closed-loop
Multi-agent
simulation
Realistic
images
Real-world
roadgraph
Safety
test cases
Multi-view
images
LiDAR
point cloud
Efficient
rendering
DATA. CitySim robinson2009citysim / NGSIM usdot2016ngsim
Bench2Drive jia2024bench2drive
nuPlan nuplan / Navsim Dauner2024NEURIPS
nuScenes caesar2020nuscenes / Waymo dataset sun2020scalability
GEN. MagicDrive gao2023magicdrive / DriveDreamer wang2023drivedreamer
SimGen zhou2024simgen
W.M. KiGRAS zhao2024kigras / SMART wu2024smart
MUVO bogdoll2023muvo
Vista gao2024vista / GAIA-1 hu2023gaia
SIM. Waymax gulino2024waymax
SUMO krajzewicz2012recent / LimSim wenl2023limsim
CARLA dosovitskiy2017carla
STRIVE rempe2022generating
MetaDrive li2022metadrive
Unisim yang2023unisim / OAsim yan2024oasim
NeuroNCAP ljungbergh2024neuroncap
DriveArena yang2024drivearena
HUGSIM zhou2024hugsim
RealEngine (Ours)
2 Related work
Autonomous driving. Recent research in autonomous driving have shifted from addressing individual tasks to exploring end-to-end planning, enabling the progressive execution from perception to ego-planning within a unified framework. Early works casas2021mp3 ; hu2022stp3 ; Chitta2023PAMI implement this by leveraging simplified auxiliary tasks, which limits the final performance. In contrast, UniAD hu2023planning and VAD jiang2023vad have advanced this paradigm by integrating a broader spectrum of driving tasks, achieving notable progress in planning tasks by producing explicit intermediate results. Additionally, recent approaches present sparse representations sparsead ; sparsedrive , employ lightweight diffusion-based generation liao2024diffusiondrive to efficiently facilitate end-to-end planning. For enabling these driving methods to be evaluated extensively and authentically, we provide a reliable testing platform here.
Figure 1: The working mechanism of RealEngine. It consists of three modules: a driving agent, a motion controller and a multi-model renderer. Given a traffic scene represented by multimodal data including multi-view images and LiDAR point cloud, the driving agent predicts the trajectory, according to which RealEngine updates the ego-motion state for all traffic participants. Moving to the next time step, the multimodal sensor data will be refreshed by the current ego-motion state, which is then used for the driving agent to make the next trajectory planing. We consider three driving situations: non-reactive, safety test, and multi-agent interaction.
Sensor simulation. The geometric reconstruction of extensive urban spaces, such as streets and cities, has played a pivotal role in autonomous driving sun2020scalability ; geiger2012we ; caesar2020nuscenes . Recently, Gaussian splatting kerbl3Dgaussians based methods have been introduced to model dynamic urban scenes. PVG chen2023periodic incorporates periodic vibration at each Gaussian primitive to represent static and dynamic objects uniformly. At the same time, explicitly decomposing scenes into independent entities has become common practice, as seen in works such as StreetGaussians yan2024street , DrivingGaussians zhou2024drivinggaussian , HUGS hugs and OmniRe chen2024omnire . Recently, LiDAR simulation studies have focused on using real data for improved realism. For instance, LiDARsim manivasagam2020lidarsim and PCGen li2023pcgen use multi-step, data-driven methods to simulate point clouds from real data. Additionally, works such as tao2023lidarnerf ; zhang2024nerf ; zheng2024lidar4d ; xue2024geonlf ; tao2024alignmif ; wu2024dynamic leverage NeRF mildenhall2020nerf for scene reconstruction and LiDAR simulation. Recent works jiang2025gslidar ; zhou2024lidarrt introduce Gaussian splatting kerbl3Dgaussians into the LiDAR reconstruction task, achieving improved reconstruction quality and rendering speed. However, sensor-acquired data is prone to occlusions, leading to information loss within scenes. Additionally, reconstructed dynamic objects can distort when viewed from alternative perspectives, limiting scene editability and novel view synthesis. To address these challenges, we propose modeling the foreground and background separately allowing them to be composed in a flexible manner, resulting in a comprehensive driving simulation platform.
Closed-loop simulation in realistic settings. Closed-loop simulation gulino2024waymax ; nuplan ; dosovitskiy2017carla ; li2022metadrive is crucial for the evaluation and deployment of AD planning systems, collecting sufficient driving statistics and improving the emergency response capability of driving agents. To make simulators more realistic, recent research has explored using existing real-world driving datasets. Bench2Drive jia2024bench2drive improves the CARLA benchmark by reconstructing the data format to be more aligned with the nuScenes dataset, bridging the gap between reinforcement learning planners and end-to-end planners. Navsim Dauner2024NEURIPS imitates closed-loop evaluation to adjust the nuPlan benchmark, introducing more comprehensive and practical metrics. With the emergence of novel 3D rendering and generation research, some works integrate these techniques into existing datasets, better simulating and constructing richer and more diverse scenarios. NeuroNCAP ljungbergh2024neuroncap utilizes NeRF mildenhall2020nerf to render novel surrounding views and creates collision scenarios to enhance measurement. Relying on powerful generative models and extensive driving data, DriveArena yang2024drivearena is capable to yield controllable and abundant real-world scenarios and realize closed-loop simulation. However, these simulators all fail to meet the fundamental criteria as mentioned earlier. This motivates the introduction of our RealEngine, a more reliable and powerful benchmark for assessing the real-world performance of driving agents.
3 Method
We present RealEngine, a reliable and comprehensive autonomous driving simulation platform capable of rendering realistic multimodal sensor data efficiently and supporting closed-loop simulation. It is composed of: (i) Simulator infrastructure including background scene and traffic participants as well as their compositions; (ii) Closed-loop driving simulation; (iii) Assessment of driving agents. The working mechanism of RealEngine is depicted in Figure 1.
Figure 2: Scene composition. We start with modeling the background scene based on real sensor data and obtaining the meshes of traffic participants either by extracting them from the reconstructed data or through manual creation. That makes a rich space for designing a variety of customizable traffic scenarios. To create a specific scenario, we select both the background scene and traffic participants, which can be integrated based on each participant’s spatial coordinates over time. This naturally enables the creation of highly diverse scenes to support extensive closed-loop simulation.
3.1 Simulator infrastructure
The advancement of recent generative models has been recently exploited for closed-loop driving simulation. For example, DriveArena yang2024drivearena streamlines a couple of generative functions, such as interactive traffic flows and novel scene synthesis, allowing to simulate virtual driving scenes. However, this approach suffers from several key drawbacks: (1) Need for large manually labelled training data for video generative model optimization; (2) Substantial running overhead, even without support to multimodal sensor data such as LiDAR point clouds (3) Inability to support multi-agent co-operation and interaction. (4) Dependence on a pre-defined high-definite map; (5) Low controllability on the traffic scenario including background scene and traffic participants, consequently causing frequent situational inconsistency over space and time; (6) Low spatiotemporal realism. While some of the above issues, such as realistic scene rendering, more flexible traffic scenario editing, and no need for training data collection, can be addressed by NeuroNCAP ljungbergh2024neuroncap , it narrowly focuses on safety test of driving agents. Further, it cannot well generalize to free form trajectories otherwise the novel view synthesis will degraded substantially.
To address all the problems mentioned above, we propose decoupling the background scene and foreground traffic participants by reconstructing each of them individually from the corresponding real sensor data, allowing the composition of diverse traffic scenarios and free form driving trajectories of multiple agents while maintaining high-fidelity novel view synthesis in multiple modalities (e.g., camera images and LiDAR).
3.1.1 Background scene models
To reconstruct realistic background environments, we adopt StreetGaussians yan2024street for camera images and GS-LiDAR jiang2025gslidar for LiDAR point clouds, chosen for their high rendering efficiency and cross-modal fidelity. Importantly, RealEngine maintains flexibility—alternative methods such as PVG chen2023periodic , OmniRe chen2024omnire , and LiDAR-RT zhou2024lidarrt can be seamlessly incorporated into our system as modular components.
Pose calibrations in nuPlan nuplan is relatively coarse, posing challenges for accurate scene reconstruction. A widely used conventional method for camera pose correction yan2024street ; chen2024omnire ; hess2024splatad involves learning a trainable correction matrix that adjusts the camera pose automatically during training. However, this method is only effective for small pose deviations and fails to converge on nuPlan, leading to suboptimal results. To address this issue, we propose a pose correction method based on LiDAR point cloud registration.
As the geometric information of the background scene remains fixed in the world coordinate system, we can align the ego vehicle’s poses across different frames by registering LiDAR points transformed to the world. We remove dynamic vehicles from the LiDAR data based on annotation information and filter out ground points which vary significantly across frames. All remaining point clouds are transformed into the world coordinate system and truncated within a predefined region to ensure consistency across frames. We select the central LiDAR frame ${\bm{P}}_{\mathrm{refer}}$ as the reference frame and apply a learnable correction matrix ${\bm{M}}$ to transform the LiDAR frames ${\bm{P}}$. The transformed LiDAR frames are then compared with the reference frame using the Chamfer Distance (C-D) fan2017chamfer as the loss function:
$\mathcal{L}_{\mathrm{cd}}=\mathrm{CD}({\bm{M}}{\bm{P}},{\bm{P}}_{\mathrm{refer}})$ (1)
Additionally, we leverage learnable image exposure transformations to handle cross-camera appearance variations, and utilizes video generative prior yang2024drivex ; yu2024viewcrafter to optimize the scene model across multiple trajectories in Section A.2. For further details on camera image and LiDAR point cloud reconstruction, please refer to Appendix A. By leveraging these advanced reconstruction techniques, we achieve precise multi-modal background reconstruction with high rendering efficiency, enabling flexible scene editing and realistic simulation of autonomous driving agents.
3.1.2 Traffic participants models
Traffic participants (e.g., vehicles, bicycles) are essential for realistic driving scenarios. To enable high-fidelity and diverse simulations, we curate a comprehensive set of 3D participant models. Conventional approaches yan2024street ; chen2024omnire ; zhou2024hugsim insert reconstructed Gaussian splatting primitives into the background scene according to predefined trajectories. While providing high-quality rendering from training viewpoints, they degrade significantly when rendering resolution, viewpoint, or object distance changes yan2024multi ; Yu2024MipSplatting ; sags . This degradation causes visual artifacts, especially in close-range interactions (e.g., collisions), and also introduces inconsistencies in lighting and shadows between foreground objects and the background scene.
To address these issues, RealEngine represents traffic participants using 3D meshes, ensuring consistent geometry across all viewpoints and distances. To achieve seamless integration, we employ a diffusion model to guide learning of scene-consistent lighting and shading, as further detailed in Section 3.1.3. Meshes are rendered using ray tracing to produce both RGB images and depth maps, enabling accurate occlusion reasoning between participants and scenes.
The 3D meshes are sourced from two complementary pipelines: (i) High-quality meshes manually created for key traffic objects. (ii) Meshes reconstructed from real-world datasets, including 360-degree images from CO3D reizenstein21co3d and 3DRealCar du20243drealcar . These are processed via 3D Gaussian Splatting (3DGS) for reconstruction, followed by mesh extraction for flexible use in scene composition.
3.1.3 Driving scenario composition
The combination of reconstructed background scenes and traffic participant models forms a rich design space for creating diverse and flexible driving scenarios. Scenario composition begins by selecting a background scene, into which a set of traffic participants is inserted. Each participant is assigned an initial position, orientation, and free-form trajectory, either manually specified or dynamically planned by a driving agent. These participants are then spatially registered into the background scene’s coordinate system. An overview is shown in Figure 2.
To ensure realistic integration between foreground participants and background scenes, we introduce a physically based rendering (PBR) process that optimizes environment light maps for consistent relighting and shadow casting. We adopt the Disney BRDF lighting model wang2023fegr , where the foreground color is computed as:
${\bm{C}}_{fg}({\bm{x}})=\int_{\bm{\Omega}}f_{r}({\bm{x}},{\bm{\omega}},{\bm{\omega}}_{o}){\bm{L}}_{i}({\bm{\omega}})|\langle{\bm{\omega}},{\bm{n}}\rangle|\mathrm{d}{\bm{\omega}}$ (2)
Here, ${\bm{x}}$ denotes the surface point where the camera ray $\omega_{o}$ intersects, ${\bm{n}}$ the surface normal, ${\bm{\omega}}$ the light direction, and ${\bm{L}}_{i}$ the incoming illumination. The BRDF $f_{r}$ models the surface reflectance.
For shadows, we assume a ground plane closely aligned to the foreground object’s base. For each camera ray ${\bm{\omega}}_{o}$, we compute the ground intersection ${\bm{x}}^{\prime}$ and its normal ${\bm{n}}^{\prime}$. Shadow rays ${\bm{\omega}}^{\prime}$ are traced across the hemisphere $\bm{\Omega}^{\prime}$ to determine occlusion, and the shadow intensity ${\bm{I}}$ is computed:
${\bm{I}}({\bm{x}}^{\prime})=\frac{\int_{\bm{\Omega}({\bm{x}}^{\prime})}{\bm{L}}_{s}({\bm{\omega}}^{\prime})|\langle{\bm{\omega}}^{\prime},{\bm{n}}^{\prime}\rangle|\mathrm{d}{\bm{\omega}}^{\prime}}{\int_{\bm{\Omega}^{\prime}}{\bm{L}}_{s}({\bm{\omega}}^{\prime})|\langle{\bm{\omega}}^{\prime},{\bm{n}}^{\prime}\rangle|\mathrm{d}{\bm{\omega}}^{\prime}}$ (3)
where ${\bm{L}}_{s}$ is the shadow environment map, and $\bm{\Omega}({\bm{x}}^{\prime})$ represents the solid angle of unoccluded light rays at point ${\bm{x}}^{\prime}$ with respect to the foreground mesh.
The final composed image is:
${\bm{C}}={\bm{C}}_{bg}\cdot{\bm{I}}\cdot(1-{\bm{M}}_{fg})+{\bm{C}}_{fg}\cdot{\bm{M}}_{fg}$ (4)
where ${\bm{C}}_{bg}$ is the background image, ${\bm{C}}_{fg}$ and ${\bm{M}}_{fg}$ are the rendered foreground RGB and mask, respectively. We optimize $\{{\bm{L}}_{i},{\bm{L}}_{s}\}$ using StableDiffusion sd with SDS loss poole2022dreamfusion , ensuring photorealistic blending.
3.2 Closed-loop driving simulation
Built upon this flexible scenario composition and novel view synthesis-enabled reconstruction, RealEngine supports a wide range of driving scenarios. We demonstrate three representative categories:
Non-reactive Simulation. Evaluates closed-loop planning where other participants follow fixed pre-recorded trajectories (Figure 1 (a)). The agent perceives rendered multi-modal sensor data, plans trajectories, and executes motion via a linear quadratic regulator (LQR), completing the closed-loop cycle.
Safety Test Simulation. Evaluates agent reactions to inserted participants exhibiting hazardous behaviors (e.g., sudden lane changes or blocked intersections), testing safety-critical skills (Figure 1 (b)).
Multi-agent Interaction Simulation. Simulates cooperative and adversarial interactions where multiple agents, each running independent planning, simultaneously navigate the scene (Figure 1 (c)).
3.3 Assessment of driving agents
Driving agent assessment aims to evaluate the agent’s trajectory planning performance across both space and time, considering safety, progress, and interaction quality within diverse traffic scenarios. In traditional open-loop evaluation, the Predictive Driver Model Score (PDMS) Dauner2024NEURIPS measures trajectory quality at individual time steps. However, open-loop PDMS does not capture how planning decisions evolve over time, nor does it account for interactions with dynamic participants in the scene.
To address these issues, RealEngine introduces a closed-loop extension of PDMS, which evaluates the agent’s performance over a continuous sequence of future steps. Specifically, starting at time step $t$, the agent plans a sequence of future positions over $N$ steps, continuously interacting with the environment and surrounding participants.
At each step, the agent receives updated sensor observations rendered from the new position, which are used to generate the next planned trajectory, creating a fully closed-loop simulation. During this process, other traffic participants follow either their pre-recorded reference trajectories or react through independent planning processes, depending on the scenario type.
PVG chen2023periodic OmniRe chen2024omnire
Ground truth
StreetGaussians yan2024street RealEngine (Ours)
(a) Comparison of camera images reconstruction (b) Relighting
Figure 3: (a) Compared to state-of-the-art reconstruction methods, we achieve superior camera image reconstruction in the nuPlan nuplan benchmark. Additionally, (b) our foreground relighting technique enables seamless integration of inserted participants with the background scene.
(a) ST-P3 hu2022stp3
(b) DiffusionDrive liao2024diffusiondrive
Figure 4: Non-reactive simulation. The driving agent’s planned trajectory at each step is visualized in bird’s-eye view (blue) and the front-view camera (green). During closed-loop simulation, ST-P3 hu2022stp3 exhibits inconsistencies in consecutive frame planning, leading to error accumulation and causing the vehicle to navigate into invalid regions. In contrast, DiffusionDrive liao2024diffusiondrive maintains more consistent planning across consecutive frames, resulting in a higher PDM Score.
The final closed-loop PDMS over the horizon $t$ to $t+N$ is formulated as:
$\mathrm{PDMS}_{t:t+N}=\underbrace{\left(\prod_{m\in\{\mathrm{NC},\mathrm{DAC}\}}\text{score}_{m}\right)}_{\text{penalty terms}}\times\underbrace{\left(\frac{\sum_{w\in\{\mathrm{EP},\mathrm{TTC},\mathrm{C}\}}\mathrm{weight}_{w}\cdot\text{score}_{w}}{\sum_{w\in\{\mathrm{EP},\mathrm{TTC},\mathrm{C}\}}\mathrm{weight}_{w}}\right)}_{\text{weighted average rewards}}.$ (5)
The penalty terms include: (1) NC (No Collision): Whether the agent avoids collisions with other vehicles, pedestrians, or cyclists. (2) DAC (Drivable Area Compliance): Whether the agent stays within the valid drivable area (lanes, intersections, etc.).
The weighted average reward terms reflect the overall driving quality across three dimensions: (1) EP (Ego Progress): How effectively the agent advances toward the goal within the given time horizon. (2) TTC (Time to Collision): How much time the agent maintains between itself and potential collision risks. (3) C (Comfort): Evaluating smoothness of the planned trajectory, including acceleration and jerk. This combined metric captures both safety-critical behaviors (penalties) and desirable driving qualities (rewards), providing a comprehensive assessment of the agent’s closed-loop performance in diverse and dynamic traffic scenarios.
4 Experiments
Datasets. We use CO3D reizenstein21co3d and 3DRealCar du20243drealcar to reconstruct diverse foreground traffic participants, followed by mesh extraction for rendering and scene composition. Additionally, we include high-quality meshes from Sketchfab Sketchfab for further diversity in foreground insertion and relighting.
For background scene reconstruction and scene editing, we leverage Navsim Dauner2024NEURIPS , which is derived from OpenScene openscene2023 , a simplified version of nuPlan nuplan . We select 14 diverse sequences from Navsim for scene editing and driving agent evaluation, and design 14, 21, and 14 test cases for the non-reactive, safety, and multi-agent interaction simulation, respectively. To ensure high-quality sensor simulation, we retrieve high-frequency images and LiDAR point clouds from the corresponding nuPlan sequences for background scene reconstruction. Using an NVIDIA RTX A6000, the rendering frame rate reaches 30Hz for camera images and 15Hz for LiDAR data.
Autonomous driving models. We evaluate four end-to-end driving models: ST-P3 hu2022stp3 , VAD jiang2023vad , TransFuser Chitta2023PAMI , and DiffusionDrive liao2024diffusiondrive , with ST-P3 and VAD reimplemented on nuPlan nuplan for consistency. As a baseline, we also test Navsim’s constant velocity model for comparison.
Table 2: Non-reactive simulation. We show the no at-fault collision (NC), drivable area compliance (DAC), time-to-collision (TTC), comfort (Comf.), and ego progress (EP) subscores, and the PDM Score (PDMS), as percentages.
Method Loop Ego stat. Image LiDAR NC $\uparrow$ DAC $\uparrow$ TTC $\uparrow$ Comf. $\uparrow$ EP $\uparrow$ PDMS $\uparrow$
Constant velocity Dauner2024NEURIPS 92.9 64.3 85.7 100 29.4 46.8
ST-P3 hu2022stp3 Open-loop 92.9 71.4 92.9 100 46.2 59.6
VAD jiang2023vad 92.9 85.7 92.9 100 48.5 66.1
TransFuser Chitta2023PAMI 92.9 85.7 92.9 100 55.9 69.1
DiffusionDrive liao2024diffusiondrive 92.9 85.7 92.9 100 56.7 69.5
ST-P3 hu2022stp3 Closed-loop 100 64.3 85.7 100 35.6 47.5
VAD jiang2023vad 85.7 78.6 78.6 100 34.3 53.0
TransFuser Chitta2023PAMI 92.9 71.4 85.7 100 46.0 57.9
DiffusionDrive liao2024diffusiondrive 92.9 71.4 92.9 100 47.1 61.3
Human 100 100 92.9 100 68.3 83.8
(a) Stationary          (ST-P3 hu2022stp3 ) (b) Sudden brake           (VAD jiang2023vad ) (c) Intersection encounter (TransFuser Chitta2023PAMI ) (d) Cut in (DiffusionDrive liao2024diffusiondrive )
Figure 5: Safety test simulation. The driving agent’s planned trajectory at each step is visualized in bird’s-eye view (blue) and the front-view camera (green). The driving agent is navigating in our designed safety-critical scenarios. The agent may (a)(b) successfully avoid a collision, (c) exhibit no reaction, or (d) attempt to decelerate to prevent a collision but ultimately fail.
4.1 Diving scenario quality
We compared our optimized reconstruction results (Appendix A) with the state-of-the-art methods yan2024street ; chen2024omnire ; chen2023periodic on 6 Navsim Dauner2024NEURIPS sequences, as shown in Figure 3 (a) and Table 3, and observed a notable improvement in the reconstruction performance on nuPlan. Meanwhile, our foreground insertion blends more harmoniously with the background and can be recognized by perception models borse2021inverseform ; xie2021segformer ; ren2024grounded , as shown in Figure 3 (b). This enables the driving agents to correctly recognize the inserted foreground participants and respond accordingly.
Table 3: Comparison with state-of-the-art camera images reconstruction methods in the nuPlan benchmark.
[…]
StreetGaussians yan2024street 26.02 0.798 0.179
Omnire chen2024omnire 26.89 0.837 0.165
PVG chen2023periodic 28.32 0.854 0.176
RealEngine(Ours) 29.67 0.897 0.093
(a) Agent 1 (DiffusionDrive liao2024diffusiondrive )
(b) Agent 2 (DiffusionDrive liao2024diffusiondrive )
Figure 6: Multi-agent interaction simulation. The trajectory annotations are consistent with Figure 5. We set two agents (DiffusionDrive liao2024diffusiondrive ) to plan trajectories simultaneously and independently in the showed turning scenario. The first agent avoids collision by increasing its turning radius, while the second agent decelerates to allow safe passage. Once the vehicles have passed each other, the second agent resumes its original speed.
Table 4: Safety test and multi-agent simulation. The notations are consistent with the non-reactive simulation Table 2 above.
Method Simulation Ego stat. Image LiDAR NC $\uparrow$ DAC $\uparrow$ TTC $\uparrow$ Comf. $\uparrow$ EP $\uparrow$ PDMS $\uparrow$
Constant velocity Dauner2024NEURIPS Safety test 47.6 71.4 38.1 100 36.7 36.3
ST-P3 hu2022stp3 47.6 100 42.9 100 44.7 44.4
VAD jiang2023vad 47.6 95.2 28.6 100 41.2 37.0
TransFuser Chitta2023PAMI 47.6 100 38.1 100 44.1 42.2
DiffusionDrive liao2024diffusiondrive 57.1 100 52.4 100 54.0 53.8
Constant velocity Dauner2024NEURIPS Multi-agent 42.8 60.7 39.3 100 27.8 27.4
ST-P3 hu2022stp3 53.6 96.4 50.0 100 44.6 46.3
VAD jiang2023vad 32.1 71.4 32.1 100 27.7 28.8
TransFuser Chitta2023PAMI 60.7 96.4 53.6 100 54.3 55.0
DiffusionDrive liao2024diffusiondrive 57.1 96.4 50.0 100 51.7 51.9
4.2 Closed-loop simulation
Non-reactive simulation. In the non-reactive setting, we conduct open-loop and closed-loop simulations on the driving agents and compare their performance, as shown in Table 2. Compared to directly predicting multiple frame trajectories in open-loop simulation, the PDM Score of the continuously predicted trajectories by the driving agent in closed-loop simulation is reduced to varying degrees. In particular, models tend to exhibit inconsistency in the adjacent predicted trajectories. This inconsistency leads to cumulative errors, causing the vehicle to navigate to unreasonable areas and resulting in a decrease in the DAC (drivable area compliance) metric, as shown in Figure 4.
Safety test simulation. For safety test simulation, we designed three test cases (simple, moderate, and challenging) tailored to each specific scenario and conducted closed-loop simulations to compute the PDM Score. As shown in Figure 5, in simple and moderate scenarios, such as a stationary vehicle blocking the lane or a sudden brake, driving agents can reasonably maneuver to avoid collisions. However, in more challenging scenarios, such as aggressive lane changes or intersection encounters, driving agents may exhibit some reaction but fail to avoid collisions, or in some cases, show minimal response to an imminent collision, leading to a significant drop in the PDM Score, as shown in Table 4.
Multi-agent interaction simulation. In the multi-agent interaction simulation, each vehicle is assigned to an instance of a certain driving agent model for independent and simultaneous control, with only the initial speed and high-level driving commands (right, left, straight, or unknown) specified. For each specific scenario, we design both simple and challenging test cases. At each time step, we render the sensor data for all driving agents to plan trajectories and conduct closed-loop simulations to compute the PDM scores for all instances and average the results, as presented in Table 4. Additionally, Figure 6 illustrates a turning scenario involving two DiffusionDrive liao2024diffusiondrive instances, where both agents successfully avoid collisions by adopting different yet reasonable strategies.
5 Conclusion
This paper introduces a novel driving simulation platform, RealEngine, capable of rendering realistic multimodal sensor data efficiently and perform closed-loop simulation in highly diverse traffic scenarios. RealEngine employs a foreground-background separate reconstruction and composition rendering approach, enabling convenient and more flexible control and editing of scenes. This allows for realistic interaction and continuous feedback between the driving agent and the simulation platform, while also facilitating multi-agent interactions within the simulation. Based on this, we simulate three driving categories: non-reactive, safety test, and multi-agent interaction, to establish a reliable and comprehensive benchmark for evaluating the real-world performance of driving agents.
Appendix A Urban scene reconstruction
A.1 NuPlan pose correction
We propose a pose correction method based on LiDAR point cloud registration in Section 3.1.1. As shown in Figure 7, point cloud registration effectively corrects pose calibration errors. This leads to improved scene reconstruction quality, which is reflected in the increased PSNR, as presented in Table 5.
(a) Origin (b) After pose correction
Figure 7: Pose correction in nuPlan. The yellow and red points originate from LiDAR point clouds of different frames. (a) Due to the coarse pose calibration in nuPlan, LiDAR point clouds from different frames are misaligned in the world coordinate system, posing challenges for high-quality reconstruction. (b) After our pose correction, LiDAR point clouds are properly aligned, leading to improved camera images reconstruction quality.
Table 5: Ablation study on camera images reconstruction in the nuPlan benchmark.
[…]
w/o $\mathcal{L}_{\text{cd}}$ 26.02 0.798 0.179
w/o undistortion 25.68 0.809 0.182
w/o color correction 28.05 0.853 0.132
RealEngine(Ours) 29.67 0.897 0.093
A.2 Camera images reconstruction
Figure 8: Distortion and color inconsistency in the nuPlan benchmark. (a) The camera images in nuPlan exhibit significant barrel distortion, which poses challenges for Gaussian splatting based reconstruction. To address this, we apply distortion correction to the images. (b) Additionally, different cameras have varying exposure levels. To prevent color ambiguity for the same Gaussian primitive across different cameras, we learn an exposure transformation for each camera separately.
We use StreetGaussians yan2024street to reconstruct the scene and perform additional processing on the nuPlan dataset. StreetGaussians yan2024street uses manual pose annotation of dynamic objects to distinguish between static background and moving objects. Dynamic objects are reconstructed in their respective centered canonical spaces and then placed into the background scene space during the rendering process based on the known poses.
As shown in Figure 8 (a), the camera images in nuPlan exhibit significant barrel distortion. Since 3DGS relies on image-based rendering and does not render rays corresponding to each pixel as NeRF does, such distortion adversely affects the reconstruction of scene geometry. To mitigate this issue, we applied a distortion correction to the images. Furthermore, due to varying exposure levels across the cameras in the nuPlan dataset, the same object exhibits substantial color differences when viewed from different cameras, as shown in Figure 8 (b), leading to ambiguities in the color information of Gaussian primitives. To address this, we introduced a learnable affine transformation $\{{\bm{A}}_{i},{\bm{t}}_{i}\}$ for each camera $i$ to individually calibrate the image colors:
$\tilde{{\bm{C}}}_{i}={\bm{A}}_{i}{\bm{C}}_{i}+{\bm{t}}_{i}$ (6)
where $i$ represents the camera index, ${\bm{C}}_{i}$ denotes the original rendered RGB map, and $\tilde{{\bm{C}}}_{i}$ refers to the RGB map after exposure transformation. These processing methods improve the quality of scene reconstruction, as shown in the ablation study in Table 5.
We compared our optimized reconstruction results with the state-of-the-art methods yan2024street ; chen2024omnire ; chen2023periodic on 6 Navsim Dauner2024NEURIPS sequences, as shown in Figure 3 and Table 3, and observed a notable improvement in the reconstruction performance on nuPlan.
Although existing urban reconstruction methods perform excellently in synthesizing novel viewpoints for recorded trajectories, they face challenges when handling new trajectories in closed-loop simulations due to the limited viewpoints of driving videos and the vastness of the driving environment. To address this challenge, we employ DriveX yang2024drivex to enhance the reconstructed scene, which utilizes video generative prior yu2024viewcrafter to optimize the scene model across multiple trajectories. As shown in Figure 9, our final scene model is capable of generating high-fidelity virtual driving environments beyond the recorded trajectories, enabling free-form trajectory driving simulations.
Figure 9: Lane change reconstruction. We render the camera images by shifting the driving perspective 3 meters to the left and right. Due to the limited viewpoint of driving videos and the vastness of the driving environment, large translations from driving perspectives lead to a significant decline in reconstruction quality. To address this, we use DriveX yang2024drivex , which leverages video generative prior to optimize the scene Gaussian primitives, resulting in improved reconstruction quality even with large viewpoint translations.
A.3 LiDAR reconstruction
Table 6: Evaluation of LiDAR reconstruction performance in the nuPlan benchmark.
Point Cloud Depth Intensity
[…]
GS-LiDAR jiang2025gslidar 0.315 6.930 21.25 0.0903 20.88
We use GS-LiDAR jiang2025gslidar to reconstruct the LiDAR point clouds. GS-LiDAR jiang2025gslidar introduces a novel panoramic rendering technique with explicit ray-splat intersection, guided by panoramic LiDAR supervision. This method employs 2D Gaussian primitives huang20242d with periodic vibration characteristics chen2023periodic , allowing for precise geometric reconstruction of both static and dynamic elements in driving scenes.
The LiDAR point clouds in nuPlan nuplan is generated by merging point clouds from five LiDAR sensors, with non-uniform angular distributions across different scan lines. However, GS-LiDAR assumes that the LiDAR point cloud originate from a single laser scanner with uniformly spaced scan lines. To address this discrepancy, we reproject the merged point cloud onto the top LiDAR sensor to obtain the range map, selecting the point with the smaller depth when multiple points project to the same pixel. Additionally, we discard the non-uniform scan lines at both ends while retaining the central region, where the scan lines are evenly distributed.
As shown in Figure 10, this processing method results in some loss of LiDAR information. However, since the driving agent Chitta2023PAMI ; liao2024diffusiondrive converts the LiDAR point cloud into a 2-bin histogram over a 2D BEV grid with relatively low resolution, the missing information has a minimal impact on the driving agent. We evaluate the reconstruction performance on the same 6 sequences as the camera image reconstruction in Table 6.
Figure 10: LiDAR reconstruction. Although reprojection may lead to some loss of LiDAR information, its impact on the low-resolution histogram used by the agent is minimal. Meanwhile, GS-LiDAR jiang2025gslidar achieves high-quality reconstruction of the reprojected LiDAR data.
Appendix B Consistency of driving behaviors
We replicate the same real-world evaluation scenarios from the Navsim Dauner2024NEURIPS benchmark in the non-reactive open-loop setting in Table 2, and evaluate autonomous driving models Chitta2023PAMI ; liao2024diffusiondrive trained solely on real data in both the real open-loop environment and its simulated counterpart in RealEngine, comparing the performance differences.
We use the the Predictive Driver Model Score (PDMS) Dauner2024NEURIPS between trajectories as a measure of similarity, and conduct evaluations only on sequences where the PDMS is greater than zero. The gap is defined as follows:
$\mathrm{gap}=\frac{\left|\mathrm{PDMS}_{\mathrm{real}}-\mathrm{PDMS}_{\mathrm{sim}}\right|}{\max(\mathrm{PDMS}_{\mathrm{real}},\mathrm{PDMS}_{\mathrm{sim}})}$ (7)
Specifically, $\mathrm{PDMS}_{\mathrm{real}}$ refers to the PDMS of the trajectory planned based on real sensor inputs, whereas $\mathrm{PDMS}_{\mathrm{sim}}$ corresponds to that planned using sensor data simulated by RealEngine. To evaluate the realism of camera and LiDAR simulation independently, we test three settings: (i) simulating only the camera ($\mathrm{PDMS}_{\mathrm{cam}}$), (ii) simulating only the LiDAR ($\mathrm{PDMS}_{\mathrm{lidar}}$), and (iii) simulating both camera and LiDAR ($\mathrm{PDMS}_{\mathrm{both}}$), as shown in Table 7. To better understand the realism of RealEngine, we also visualize the trajectories planned by DiffusionDrive liao2024diffusiondrive , as shown in Figure 11.
Table 7: Consistency of driving behaviors. We test Transfuser Chitta2023PAMI and DiffusionDrive liao2024diffusiondrive , which take camera and LiDAR inputs, under the three settings described in Appendix B.
Method Image LiDAR $\mathrm{PDMS}_{\mathrm{real}}\uparrow$ $\mathrm{PDMS}_{\mathrm{cam}}\uparrow$ $\mathrm{gap}_{\mathrm{cam}}\downarrow$ $\mathrm{PDMS}_{\mathrm{lidar}}\uparrow$ $\mathrm{gap}_{\mathrm{lidar}}\downarrow$ $\mathrm{PDMS}_{\mathrm{both}}\uparrow$ $\mathrm{gap}_{\mathrm{both}}\downarrow$
TransFuser Chitta2023PAMI 81.18 80.99 0.77% 80.78 0.55% 80.56 0.98%
DiffusionDrive liao2024diffusiondrive 80.75 81.07 0.89% 80.85 0.42% 81.07 1.07%
Figure 11: Consistency of driving behaviors. We test DiffusionDrive liao2024diffusiondrive under the same settings in Table 7. The close alignment between trajectories in the real and simulated environments demonstrates the high fidelity of our RealEngine.
The results show that the planned trajectory gap between the real-world environment and the counterpart simulated by RealEngine remains small (approximately 1%), which highlights the realism of our simulator and its effectiveness in supporting closed-loop evaluation of autonomous driving models.
Appendix C Limitations
We still have a few limitations for future improvement:
Firstly, our current foreground meshes rely on existing datasets reizenstein21co3d ; du20243drealcar and 3D asset libraries, which limits the diversity of the foreground. We plan to incorporate generative models to generate foreground objects, thereby reducing manual effort, enhancing the richness of the scenes, and improving the generalizability of the simulation platform.
Secondly, although we have relighting models, there are still noticeable differences between the inserted traffic participants and the background scene. To address this issue, we plan to continuously track and integrate state-of-the-art advancements in foreground insertion to ensure seamless blending of traffic participants with the background.
Thirdly, we employ a simplified vehicle model that does not account for aerodynamic effects, suspension, or tire characteristics. Additionally, we overlook substantial variations in road conditions, such as changes in friction coefficients (e.g., wet or icy surfaces), gradients, and irregularities.